Suppose a firm produces according to the production function Q = 2L0.6K0.2, and faces wage rate ₵10, a rental cost of capital ₵5, and sells output at a price of ₵20.

a. Obtain and expression for the factor demand functions.

b. Compute the profit-maximizing factor demands for capital and labour.


1
Expert's answer
2021-10-18T11:27:28-0400

"Q=2L^{0.6}K^{0.2} \\

w = 10 \\

r = 5 \\

p = 20"

a. Input demand for labor L = f(w,r,p)

Input demand for capital K=f(w,r,p)

Profit

"\u03a0 = p(2L^{0.6}K^{0.2}) -wL -rK"

At maximum profit and low cost (minimum) ratio of capital to labor is

"\frac{K}{L} = \frac{(w/p)^5}{(r/p)^{5/3}} = \frac{(1/2)^5}{(1/2)^{10/3}} \\

= (\frac{1}{2})^{5/3} = 0.315"

Implies more capital required for one unit increase in wage rate.

b. Profit "\u03a0 = pQ-wL-rK"

"\u03a0 = p(2L^{0.6}K^{0.2}) -wL -rK"

For maximim profit

"\frac{dΠ}{dL} = \frac{dΠ}{dL} =0 \\

\frac{d}{dL}(p(2L^{0.6}K^{0.2}))-wL-rK = 0 \\

pK^{0.2} -w= 0 \\

K=(\frac{w}{p})^5 = (\frac{10}{20})^5 = (\frac{1}{2})^5 = 0.03125 \\

\frac{d}{dK}(p(2L^{0.6}K^{0.2}))-wL-rK =0 \\

pL^{0.6} -r=0 \\

L=(\frac{r}{p})^{5/3} = (\frac{5}{20})^{5/3} \\

L = (\frac{1}{4})^{5/3} = 0.099"